OpenStudy (anonymous):

A stone is dropped from a height of 7.70m. How fast is it going when it reaches the ground?

OpenStudy (anonymous):

Same thing. The potential energy at top will be equal to kinetic energy at bottom.

OpenStudy (anonymous):

i think i got it... im not sure though

OpenStudy (anonymous):

It doesn't give us a mass or velocity

OpenStudy (anonymous):

velocity you have to calculate.. mgh = 1/2 mv^2 v = sqrt(2gh) See the velocity is independent of mass :)

OpenStudy (anonymous):

i got 12.2849501423 and it was correct! thank you so much :) btw im a girl

OpenStudy (anonymous):

Oh, Girl you are sorry :D

OpenStudy (anonymous):

wanna help me again?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A 1600 kg car on a flat ground is moving 6.25 m/s. Its engine creates a 1150N forward force as the car moves 45.8m. What is its final KE?